# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 3.3 Simple examples
In the following we present examples of exogenous information structures for the simple examples from the previous section, subsection 2.3.
@@
    1. (a)
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\mathcal{P}(\Omega)$: only at the second move, it becomes known which scenario is realised, irrespective of which one is the second move;
    2. (b)
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\mathcal{P}(\Omega)$, $\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\{\Omega,\emptyset\}$: ${\bm{\mathsf{x}}}_{1}$ is the only move at that the realised scenario is revealed; an agent with this exogenous information may have interest in choosing (if possible) ${\bm{\mathsf{x}}}_{1}$ rather than ${\bm{\mathsf{x}}}_{2}$ in order to learn, modelling the trade-off exploration vs. exploitation; that way, problems with partial information and adaptive control can be modelled;
    3. (c)
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\mathcal{P}(\Omega)$, $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\{\Omega,\emptyset\}$: analogous to the preceding situation;
@@
    $\mathscr{F}_{\bm{\mathsf{x}}}=\mathcal{P}(\Omega)$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$: at all moves, the realised scenario is known.
@@
For the simple stochastic decision forest $(F,\pi,{\bm{\mathsf{X}}})$ from subsection 2.3 there are exactly five exogenous information structures, and these are given by the families $\mathscr{F}$ considered just above.
@@
For the stochastic decision forest $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ from subsection 2.3 there are exactly three exogenous information structures, and these are given by the families $\mathscr{F}^{\prime}$ considered just above.
@@ 3.4 Action path stochastic decision forests
In this subsection, we discuss examples of exogenous information structures for the action path sdf $(F,\pi,{\bm{\mathsf{X}}})$ induced by $W\subseteq\Omega\times\mathbb{A}^{\mathbb{T}}$ from subsection 2.4, defined over an exogenous scenario space $(\Omega,\mathscr{A})$. First, we observe a necessary condition on these: when evaluated along maximal chains of random moves in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$, we obtain a filtration in the sense of classical probability theory. More precisely:
@@
Let $\mathscr{F}$ be an exogenous information structure for the action path sdf induced by a set $W$ as in subsection 2.4, and let $f\in\mathbb{A}^{\mathbb{T}}$. Let $\mathbb{T}_{f}:=\{t\in\mathbb{T}\mid D_{t,f}\neq\emptyset\}$.
@@
Now we turn to sufficient conditions, or more precisely, to the construction of exogenous information structures for the action path sdf. Let $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ be a filtration on $(\Omega,\mathscr{A})$. For example, $\mathscr{G}$ could be generated by some stochastic process, possibly augmented with nullsets with respect to some set of probability measures on $(\Omega,\mathscr{A})$.
@@
    The basic setting in stochastic control of exogenous noise revealed over time independently of the agents’ behaviour (see, e.g., [34] corresponds to letting, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$,
@@
    and $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$.
@@
    We can also consider a more general case that allows for exogenous information depending on previous behaviour – which can serve as a basis for describing problems with partial information involving stochastic filtering, the trade-off exploration vs. exploitation, and adaptive control (see, e.g., [8, 14, 15]). Again, let $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ be a filtration on $(\Omega,\mathscr{A})$. Let $(Y_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ be a family of random variables with values in some metric space $\mathbb{B}$. Let, for any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$:
@@
    Again, let $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$.
In particular, these two cases can be used to model “open loop” and “closed loop” controls, respectively, covering the definitions in [13, chapter 2], for example. However, as will be discussed in the section 4 about adapted choices, the crux lies in that local decisions not only depend on the filtration-like object representing exogenous information at the current random move, but also on the current random move itself. Hence, when comparing a counterfactual random move to a reference random move revealing the same exogenous information, the same strategy may demand a different choice, because the random move is different.
@@
for all $t\in\mathbb{T}$ and $f\in C(\mathbb{R})$, and bounded continuous $\tilde{b}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and a suitable $\mathscr{G}$-adapted stochastic integrator $\tilde{Z}$, say Brownian motion. Then let $(Y_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ solve the stochastic differential equation
@@
Consider the action path sdf induced by a set $W$, and the family $\mathscr{F}$ from above, in its general version as in point (2). Then, $\mathscr{F}$ defines an exogenous information structure on that sdf.
@@ 4.1 Choice under uncertainty and adapted processes
In the refined partitions approach classical in sequential decision theory and developed in high generality in [3], a choice consists in selecting a member of a given partition of the set of outcomes that are still possible according to the information the given agent has at the current move. The partition members can be seen as “local outcomes”, pertaining to the current information. This is directly in line with the traditional approach of decision theory in the spirit of Savage ([38]): a strategy will then be nothing else than a Savage act selecting a local outcome at each move (alias “state” in Savage’s terms) in such a way that the set of possible outcomes is progressively reduced to a singleton.
@@
In probabilistic models based on filtrations describing exogenous information on an abstract measurable space $(\Omega,\mathscr{A})$, local choices are modelled differently. If $\mathscr{F}=(\mathscr{F}_{t})_{t\in\mathbb{T}}$ is a filtration on $(\Omega,\mathscr{A})$ with some totally ordered time index set $\mathbb{T}$, then the choice of an agent $i$ equipped with that information and capable of action described by a metric space $\mathbb{A}^{i}$ at a time $t$ is typically modelled by an $\mathscr{F}_{t}$-measurable function $g_{t}\colon\Omega\to\mathbb{A}^{i}$, modulo regularity conditions in $t$ (leading to predictable processes, for instance), see, for instance, [34, 14]. $\mathscr{F}_{t}$-measurability of $g_{t}$ means that for any measurable set of actions $A_{t}^{i}\subseteq\mathbb{A}^{i}$, the event that $i$ chooses an action in $A_{t}^{i}$ is measurable to agent $i$ at time $t$, that is, $(g_{t})^{-1}(A_{t}^{i})\in\mathscr{F}_{t}$. Actually, $A_{t}^{i}$ can be understood as a set-valued reference choice and $(g_{t})^{-1}(A_{t}^{i})$ is just the event that both choices, $g_{t}$ and $A^{i}_{t}$, are compatible.
@@
In the discrete case, this approach can be subsumed under the refined partitions framework going back to [45] and presented in high generality in [5]. Indeed, if $\mathscr{A}$ is generated by a countable partition, then so is $\mathscr{F}_{t}$ for each $t\in\mathbb{T}$, and $\mathscr{F}_{t}$-measurability is another way of expressing the requirement that $g_{t}$ is constant on each member of the partition generating $\mathscr{F}_{t}$. Recall from subsection 3.1 that each realised partition member is thought of as the agent’s information set regarding the “nature” agent’s past choices. In that sense, $g_{t}$ corresponds to a map from moves at time $t$ to action such that at moves from the same information set the same action is selected. A family of such maps $g_{t}$, ranging over all times $t$, also called adapted process in the language of stochastic processes, then defines a complete contingent plan of action for this agent $i$ provided $i$ has only exogenous information.
@@
On general measurable spaces, however, the $\mathscr{F}_{t}$-measurability cannot be rephrased like this in terms of partitions. Moreover, agents may have endogenous information about the own or other agents’ past behaviour and condition their own behaviour on this information, that is, they may condition on what they know about their current random move. This suggests that one may adapt the more general measure-theoretic concept of measurable functions to stochastic decision forests as regards exogenous information about the horizontal $\Omega$ axis, while maintaining the refined partitions logic along the vertical tree axis in order to rigorously explain the interactive decision making in extensive form. The aim of this section is to introduce such a concept, bringing together the measure-theoretic probabilistic and the refined partitions based tree-like approaches to information. [^13] In this section we do however not make formally precise which axioms the set of choices an agent is equipped with should satisfy in order to consistently define a stochastic extensive form following the refined partitions paradigm and how exactly information sets can be modelled, nor do we make precise how and when this general stochastic approach can be represented using a “nature” agent: this is the aim of the follow-up paper.
@@ 4.2 Choices in stochastic decision forests
A basic principle in extensive form theory is that at any “move” it is “known” to decision makers whether a given “choice” is available to them or not. One of the important facts the refined partitions approach formalised in [3] elucidates, is that the availability of a choice at a given move can be completely described in terms of the underlying set-theoretic structure: A choice is available at a move iff the latter is an immediate predecessor of the former. More precisely, if $(F,\pi,{\bm{\mathsf{X}}})$ is a stochastic decision forest over an exogenous scenario space $(\Omega,\mathscr{A})$, $W=\bigcup F$ and $c\subseteq W$ is some subset (for instance, a union of nodes representing a choice), then, with
@@
For each $A\subset\Omega$, and each subset $c\subseteq W$,
$P(c\cap W_{A})=P(c)\cap F_{A}.$
As in [3], a choice is a non-empty union $c$ of nodes. For $x\in X$, $c$ is said to be available at $x$ iff $x\in P(c)$, as in [3]. The sets $P(c)$ are a model for information sets in [3] and [4, 2] and will turn out to have, in a weaker sense, a similar role in the setting of stochastic decision forests, as discussed in the second paper.
While in a discrete setting the description of availability of choices in terms of predecessors makes it relatively easy to interpret and thus adhere to the above-mentioned basic principle, this becomes less evident in the present general measure-theoretic context, because of the tension between the discreteness of endogenous choices and the general measure-theoretic description of exogenous information in terms of systems of “measurable” sets rather than using partitions. We therefore explicitly assume a locally defined structure of reference choices like the $A_{t}^{i}$ above, that can be measured by agents at this random move. Choices taken by agents must be adapted with respect to their individual exogenous information structure and that structure of reference choices. Hence, on the one hand, choices can be “discrete” in the sense of constituting partitions and thus eliminating a sufficient amount of alternatives so that progressive choosing results in a unique outcome. And on the other hand, choices can be adapted in the sense that the availability of its restriction to any reference choice is a measurable event with respect to the current exogenous information.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest over an exogenous scenario space $(\Omega,\mathscr{A})$ and let $\mathscr{F}$ be an exogenous information structure.
@@
    2. (b)
      complete iff for every random move ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, ${\bm{\mathsf{x}}}^{-1}(P(c))$ is either empty or equal to $D_{\bm{\mathsf{x}}}$.
@@
    For any random move ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, a non-redundant and complete choice $c$ is said available at ${\bm{\mathsf{x}}}$ iff ${\bm{\mathsf{x}}}^{-1}(P(c))=D_{\bm{\mathsf{x}}}$.
@@
    A reference choice structure is a family $\mathscr{C}=(\mathscr{C}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ of sets $\mathscr{C}_{\bm{\mathsf{x}}}$ of non-redundant and complete choices available at ${\bm{\mathsf{x}}}$.
@@
    Let $\mathscr{C}$ be a reference choice structure. An $\mathscr{F}$-$\mathscr{C}$-adapted choice is a non-redundant and complete choice $c$ such that for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ that $c$ is available at and all $c^{\prime}\in\mathscr{C}_{\bm{\mathsf{x}}}$:
@@
From the agents’ perspective, both for exogenous information revelation and adapted choices the relevant tree-like object is the decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$. In that respect, choices are made on that tree with respect to the exogenous information revealed along it. Regarding outcomes and outcome generation, however, the $\Omega$ dimension and thus the forest $F$ are crucial. ${\bm{\mathsf{X}}}$ builds the link between both, and exogenous information as well as adapted choices are defined with respect to ${\bm{\mathsf{X}}}$ and such as to be compatible with each other. It is more general than usual continuous time stochastic control and differential games formulations (as in [34, 14, 13]) because of the dependence on ${\bm{\mathsf{X}}}$, rather than on the more rigid notion of time. This point, among others, will be clarified in the following examples.

### v2

@@ 3.3 Simple examples
In the following we discuss all exogenous information structures admitting recall for the simple examples from the previous section, Subsection 2.3.
@@
    1. (2(a))
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\mathcal{P}\Omega$: only at the second move, it becomes known which scenario is realised, irrespective of which one is the second move;
    2. (2(b))
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\mathcal{P}\Omega$, $\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\{\Omega,\emptyset\}$: ${\bm{\mathsf{x}}}_{1}$ is the only move at that the realised scenario is revealed; an agent with this exogenous information may have interest in choosing (if possible) ${\bm{\mathsf{x}}}_{1}$ rather than ${\bm{\mathsf{x}}}_{2}$ in order to learn, modelling the trade-off exploration vs. exploitation; that way, problems with partial information and adaptive control can be modelled;
    3. (2(c))
      $\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}=\mathcal{P}\Omega$, $\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}=\{\Omega,\emptyset\}$: analogous to the preceding situation;
@@
    $\mathscr{F}_{\bm{\mathsf{x}}}=\mathcal{P}\Omega$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$: at all moves, the realised scenario is known.
@@
For the simple stochastic decision forest $(F,\pi,{\bm{\mathsf{X}}})$ from Subsection 2.3 there are exactly five exogenous information structures on ${\bm{\mathsf{X}}}$ admitting recall, and these are given by the families $\mathscr{F}$ considered just above.
@@
For the stochastic decision forest $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ from Subsection 2.3 there are exactly three exogenous information structures on ${\bm{\mathsf{X}}}^{\prime}$ admitting recall, and these are given by the families $\mathscr{F}^{\prime}$ considered just above.
We conclude this subsection with a note on the absent-minded driver story as modelled by Gilboa in [22]. Let $I=\{1,2\}$ and $\xi^{1},\xi^{2}$ real-valued random variables on $(\Omega,\mathscr{E})$ such that there is a probability measure ${\mathbb{P}}$ on it making $(\rho,\xi^{1},\xi^{2})$ independent. On the basis of the sdf $(F,\pi,{\bm{\mathsf{X}}})$ proposed in Subsection 2.3, let $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{i}}=\sigma(\xi^{i})$ for both $i\in I$, that is, the agent active at ${\bm{\mathsf{x}}}_{i}$ has no non-trivial information about $\rho$ if both agents have posterior beliefs derived from ${\mathbb{P}}$. These give trivially rise to exogenous information structures on ${\bm{\mathsf{X}}}$ admitting recall.
@@ 3.4 Action path stochastic decision forests
In this subsection, we discuss examples of exogenous information structures for action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ and the induced action path sdf $(F,\pi,{\bm{\mathsf{X}}})$ as in Subsection 2.4, defined on an exogenous scenario space $(\Omega,\mathscr{E})$. First, we observe a necessary condition on these in the case of recall: When evaluated along maximal chains of random moves in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$, we obtain a filtration in the sense of classical probability theory. More precisely:
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be the action path sdf induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$. Let $\mathscr{F}$ be an exogenous information structure on ${\bm{\mathsf{X}}}$ admitting recall, and let $f\in\mathbb{A}^{\mathbb{T}}$. Let $\mathbb{T}_{f}=\{t\in\mathbb{T}\mid D_{t,f}\neq\emptyset\}$.
@@
Now we turn to sufficient conditions, or more precisely, to the construction of exogenous information structures for the action path sdf. Let $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ be a filtration on $(\Omega,\mathscr{E})$. For example, $\mathscr{G}$ could be generated by some stochastic process, possibly augmented with nullsets with respect to some set of probability measures on $(\Omega,\mathscr{E})$.
@@
    The basic setting in stochastic control of exogenous noise revealed over time independently of the agents’ behaviour (see, e.g. [38]) corresponds to letting, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$,
@@
    and $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$. This exogenous information structure admits recall.
@@
    We can also consider a more general case that allows for exogenous information depending on previous behaviour – which can serve as a basis for describing problems with partial information involving stochastic filtering, the trade-off exploration vs. exploitation, and adaptive control (see, e.g. [8, 14, 15]). Again, let $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ be a filtration on $(\Omega,\mathscr{E})$. Let $(Y_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ be a family of random variables with values in some metric space $\mathbb{B}$. Let, for any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$:
@@
    Again, let $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$, yielding an exogenous information structure admitting recall.
  3. 3.
    Taking $\mathscr{F}_{\bm{\mathsf{x}}}=\sigma(Y_{\bm{\mathsf{x}}})$, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, can and does often provide an exogenous information structure not admitting recall.
In particular, these cases can be used to model “open loop” and “closed loop” controls, respectively, covering the definitions in [13, Chapter 2], for example. However, as will be discussed in the Subsection 4 about adapted choices, the crux lies in that local decisions not only depend on the filtration-like object representing exogenous information at the current random move, but also on the current random move itself. Hence, when comparing a counterfactual random move to a reference random move revealing the same exogenous information, the same strategy may demand a different choice, because the random move is different.
@@
for all $t\in\mathbb{T}$ and $f\in C(\mathbb{R})$, and bounded continuous $\tilde{b}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and a suitable $\mathscr{G}$-adapted stochastic integrator $\tilde{Z}$, say Brownian motion151515… with respect to a probability measure on $(\Omega,\mathscr{E})$. Then let $(Y_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ solve the stochastic differential equation
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be the action path sdf induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$. Consider the family $\mathscr{F}$ from above, in its general version as in point (2). Then, $\mathscr{F}$ defines an exogenous information structure on ${\bm{\mathsf{X}}}$ admitting recall.
@@ 4.1 Choice under uncertainty and adapted processes
In the refined partitions approach classical in sequential decision theory and developed in high generality in [3], a choice consists in selecting a member of a given partition of the set of outcomes that are still possible according to the information the given agent has at the current move. The partition members can be seen as “local consequences”, pertaining to the current information. Upon relabelling moves as “local states”, this is a direct application of Savage’s concept of acts ([45]) which assign a (local) consequence to each (local) state. If the overall problem is well-posed this happens in such a way that the set of possible outcomes (which can be seen as global consequences) is progressively reduced to a singleton. [^16] This point is discussed in the follow-up paper.
@@
In probabilistic models based on filtrations which describe exogenous information on an abstract measurable space $(\Omega,\mathscr{E})$, choices are modelled differently. If $\mathscr{F}=(\mathscr{F}_{t})_{t\in\mathbb{T}}$ is a filtration on $(\Omega,\mathscr{E})$ with some totally ordered time index set $\mathbb{T}$, then the choice of an agent $i$ equipped with that information and capable of action described by a metric space $\mathbb{A}^{i}$ at a time $t$ is typically modelled by an $\mathscr{F}_{t}$-measurable function $g_{t}\colon\Omega\to\mathbb{A}^{i}$, modulo regularity conditions in $t$ (leading to predictable processes, for instance), see, for instance, [38, 14]. $\mathscr{F}_{t}$-measurability of $g_{t}$ means that for any measurable set of actions $A_{t}^{i}\subseteq\mathbb{A}^{i}$, the event that $i$ chooses an action in $A_{t}^{i}$ is measurable to agent $i$ at time $t$, that is, $(g_{t})^{-1}(A_{t}^{i})\in\mathscr{F}_{t}$. Actually, $A_{t}^{i}$ can be understood as a set-valued reference choice and $(g_{t})^{-1}(A_{t}^{i})$ is just the event that both choices, $g_{t}$ and $A^{i}_{t}$, are compatible.
@@
In the discrete case, this approach can be subsumed under the refined partitions framework going back to [52] and presented in high generality in [5]. Indeed, if $\mathscr{E}$ is generated by a countable partition, then so is $\mathscr{F}_{t}$ for each $t\in\mathbb{T}$, and $\mathscr{F}_{t}$-measurability is another way of expressing the requirement that $g_{t}$ is constant on each member of the partition generating $\mathscr{F}_{t}$. Recall from Subsection 3.1 that each realised partition member is thought of as the agent’s information set regarding the “nature” agent’s past choices. In that sense, $g_{t}$ corresponds to a map from moves at time $t$ to action such that at moves from the same information set the same action is selected. A family of such maps $g_{t}$, ranging over all times $t$, also called adapted process in the language of stochastic processes, then defines a complete contingent plan of action for this agent $i$ provided $i$ has only exogenous information.
@@
On general measurable spaces, however, the $\mathscr{F}_{t}$-measurability cannot be rephrased like this in terms of partitions. Moreover, agents may have endogenous information about the own or other agents’ past behaviour and condition their own behaviour on this information, that is, they may condition on what they know about their current random move. This suggests that one may adapt the more general measure-theoretic concept of measurable functions to stochastic decision forests as regards exogenous information about the horizontal $\Omega$ axis, while maintaining the refined partitions logic along the vertical tree axis in order to rigorously explain the interactive decision making in extensive form. The aim of this section is to introduce such a concept, bringing together the measure-theoretic probabilistic and the refined partitions based tree-like approaches to information. [^17] In this section we do however not make formally precise which axioms the set of choices an agent is equipped with should satisfy in order to consistently define a stochastic extensive form following the refined partitions paradigm and how exactly information sets can be modelled, nor do we discuss how and when this general stochastic approach could be represented using a “nature” agent: this is the aim of the second part of the present series, see [40].
@@ 4.2 Choices in stochastic decision forests
A basic principle in extensive form theory is that at any “move” it is “known” to decision makers whether a given “choice” is available to them or not. One of the important facts the refined partitions approach formalised in [3] elucidates, is that the availability of a choice at a given move can be completely described in terms of the underlying set-theoretic structure: A choice is available at a move iff the latter is an immediate predecessor of the former. More precisely, if $(F,\pi,{\bm{\mathsf{X}}})$ is a stochastic decision forest on an exogenous scenario space $(\Omega,\mathscr{E})$, $W=\bigcup F$ and $c\subseteq W$ is some subset (for instance, a union of nodes representing a choice), then, with
@@
For each subset $E\subseteq\Omega$, and each subset $c\subseteq W$,
$P(c\cap W_{E})=P(c)\cap F_{E}.$
As in [3], a choice is a non-empty union $c$ of nodes. For $x\in X$, $c$ is said to be available at $x$ iff $x\in P(c)$, as in [3]. The sets $P(c)$ are a model for information sets in [3] and [4, 2] and will turn out to have, in a weaker sense, a similar role in the setting of stochastic decision forests, as discussed in the second paper, see [40, Subsection 1.3].
While in a discrete setting the description of availability of choices in terms of predecessors makes it relatively easy to interpret and thus adhere to the above-mentioned basic principle, this becomes less evident in the present general measure-theoretic context, because of the tension between the discreteness of choices and the general measure-theoretic description of exogenous information in terms of systems of “measurable” sets rather than using partitions. We therefore explicitly assume a locally defined structure of reference choices like the $A_{t}^{i}$ above, that can be measured by agents at this random move. Choices made by agents must be adapted with respect to their individual exogenous information structure and that structure of reference choices. Hence, on the one hand, choices can be “discrete” in the sense of constituting partitions and thus eliminating a sufficient amount of alternatives so that progressive choosing results in a unique outcome. And on the other hand, choices can be adapted in the sense that the availability of its restriction to any reference choice is a measurable event with respect to the current exogenous information.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest on an exogenous scenario space $(\Omega,\mathscr{E})$, let $\tilde{\bm{\mathsf{X}}}\subseteq{\bm{\mathsf{X}}}$, and let $\mathscr{F}$ be an exogenous information structure on $\tilde{\bm{\mathsf{X}}}$.
@@
    2. (1(b))
      $\tilde{\bm{\mathsf{X}}}$-complete iff for every random move ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}$, ${\bm{\mathsf{x}}}^{-1}(P(c))$ is either empty or equal to $D_{\bm{\mathsf{x}}}$;
    3. (1(c))
      complete iff it is ${\bm{\mathsf{X}}}$-complete.
@@
    For any random move ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, a choice $c$ is said available at ${\bm{\mathsf{x}}}$ iff ${\bm{\mathsf{x}}}^{-1}(P(c))=D_{\bm{\mathsf{x}}}$.
@@
    A reference choice structure on $\tilde{\bm{\mathsf{X}}}$ is a family $\mathscr{C}=(\mathscr{C}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}}$ of sets $\mathscr{C}_{\bm{\mathsf{x}}}$ of non-redundant and $\tilde{\bm{\mathsf{X}}}$-complete choices available at ${\bm{\mathsf{x}}}$, ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}$.
@@
    Let $\mathscr{C}$ be a reference choice structure on $\tilde{\bm{\mathsf{X}}}$. An $\mathscr{F}$-$\mathscr{C}$-adapted choice is a non-redundant and $\tilde{\bm{\mathsf{X}}}$-complete choice $c$ such that for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}$ that $c$ is available at and all $c^{\prime}\in\mathscr{C}_{\bm{\mathsf{x}}}$:
@@
From the an agent’s perspective, both for exogenous information revelation and adapted choices the relevant order (or even tree-like) structure is the partial order (or even decision tree) $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$, and more precisely some subset $\tilde{\bm{\mathsf{X}}}$ describing the moves of that agent. In that respect, choices are made on that tree with respect to the exogenous information revealed along it. Regarding outcomes and outcome generation, however, the $\Omega$ dimension and thus the forest $F$ are crucial. ${\bm{\mathsf{X}}}$ builds the link between both, and exogenous information as well as adapted choices are defined with respect to ${\bm{\mathsf{X}}}$ and such as to be compatible with each other. It is more general than usual continuous-time stochastic control and differential games formulations (as in [38, 14, 13]) because of the dependence on ${\bm{\mathsf{X}}}$, rather than on the more rigid notion of time. This point, among others, will be clarified in the following examples.